# A pair polarimeter for multi-GeV -rays

### v1

@@ Abstract
The polarization of multi-GeV photon is an experimental observable holding many promises in multiple fields. In Astrophysics, such high-energy -rays are emitted by supermassive blackholes, pulsars or heavy-mass binary systems. Accessing their polarization would reveal much about the environment and structure of these astrophysical objects [1]. The structure of much smaller objects such as hadrons would greatly benefit from the knowledge of the -rays polarization. Indeed deeply virtual Compton scattering is a process in which a virtual photon inelastically scatters off a quark confined in the proton. The polarization of the final state photon is expected to be a key observable to study linearly polarized gluons in the proton, considered as the intrinsic strong field to form the hadron. Finally the structure of vacuum itself can be probed as well by studying the polarization of multi-GeV photons. Exposed to an intense field being either electromagnetic, gravitational or even strong such as in a quark-gluon Plasma, vacuum may become birefringent which would break Lorentz invariance. The birefringence can be highlighted by the alteration of the photon polarization when passing through the vacuum [1, 2, 3]. Unfortunately such sources or physics phenomena have low brightness or cross sections. Consequently, to measure the associated linear polarization of a multi-GeV photon with a reasonable statistical accuracy, the polarimeter must have a high Figure-of-Merit.
@@ 1.2 State-of-the-art in -polarimetry via pair production
At low energy, 10-20 MeV, a Time Projection Chamber (TPC) offers a resolution good enough to resolve the lepton pair: The photon converts in the gaz and the lepton pair is seen by the ionization of the latter from the conversion vertex. A prototype of such a concept named HARPO [4] was successfully built and caracterized by CEA, CNRS and École Polytechnique. However the maximal photon energy for polarimetry is limited by the TPC spatial resolution and the conversion efficiency is low as the converter is a gaz. Now there are pair polarimeters with very thin solid converters (0.2% ) and trackers located downstream of a dipole to resolve the lepton pair [5]. However the size and the low conversion efficiency makes it unable to study polarization of at most one million uncollimated photons. The most relevant design of a pair polarimeter for high-energy -rays is presented in Eingorn et al. [6] and shown in Figure 1.
@@ 2 A revisited concept of pair polarimeter for DVCS
The design of polarimeter proposed by Eignorn et al. has greatly inspired the subsequent work. For multi-GeV photons, a first limitation comes from the spatial resolution of MSDs which is not good enough unless the length of the cell is dramatically increased. Here we suggest to use monolithic active pixel silicon sensors (MAPS). With pixels being 2525 m2, MAPS detectors achieve 5m-resolution. An example of such detector is the Muon Forward Tracker of ALICE consisting of ALPIDE sensors [7]. Another significant advantage is the thinness of the silicon layer being only 50 m with a detection efficiency above 99%. Including the support layers (Kapton, Alumimium, Carbon,…) the material budget in the active region is approximately 0.25% X/X0. Nowadays an intense R&D effort is still ongoing to reduce even more the material budget of MAPS as well as their power consumption. However the low radiation length of MAPS decreases the efficiency of the polarimeter. To compensate this loss without increasing the number of MAPS layer and consequently the cost of the apparatus, an independent converter must be added to the cell design.
Looking back to Eingorn’s design, a second limitation stems from the thinness of the converter: All the multiple scattering happens in the vicinity of the conversion vertex and the information about the conversion angles are immediately/maximally blurred. Here we suggest to fill the cell with a low-density material as displayed by Figure 2. Unlike a thin converter, the multiple scattering in a low-density material will happen all along the track instead of immediately after the conversion. As the information is recovered from the position measurements in the MAPS, a deviation of the particle is much less significant when it occurs close to the measurement planes than at the conversion vertex. The extended converter could be some high-pressure heavy gaz or low-density material such as Rohacell 31/51/71.
@@ 2.1 A realistic MAPS description
First single-particle response will be optimized to reproduce the ALPIDE characteristics presented in [7]. Then a study of cluster shapes caused by two close particles going through the MAPS plane is performed to extract as much information as possible to be used at a later stage in a Kalman filter for the polarimeter study.
@@ 2.1.1 MAPS response to single particle
where the definitions of the plasma energy and the parameter can be found in the section entitled passage of particles through matter in the PDG [8]. The width of the distribution is . Following this distribution, an energy deposit is randomly generated for each particle in the Toy Monte-Carlo. This energy is then converted into a number of ionized electron/hole pairs placed uniformly along the path of the particle through the fully depleted zone.
@@
where is the hole diffusion constant (=10.86 cm2/s). In order to reproduce as best as possible ALPIDE caracteristics presented in the article [7], the MPV and width of the Landau-Vavilov distribution had to be multiplied by 0.28: For very thin absorber such as 50m-thick silicon sensors, the MPV is known to be smaller than the results from the PDG formula. Displayed by Figure 3, both efficiency and cluster size distributions as function of the threshold are faithfully reproduced whereas the resolution is found 1m better than the measured ones.
In addition to these fundamental properties, the average cluster size as function of the particle location in a pixel are compared in Figure 4. When the particle hit the middle of a pixel, the average cluster size is close to 1. As the particle gets closer to a neighbor pixel either on the same row or column, the average cluster size increases to 2. When getting closer to two neighbor pixels by the diagonale of the pixel, L-shaped 3-pixel clusters become the main pattern which may turn into a squared 4-pixel cluster when the particle hit the corner of a pixel.
@@ 2.1.2 MAPS response to two close particles
A single particle is first shot at a random position in middle pixel of a 99-pixel matrix. Then, at a random distance of at most 100 m from the first particle, a second particle is generated. The threshold is set to 90 electrons to consider a pixel being fired. The matrix is then analyzed to determine the number of clusters and their relative position resulting from the convolution between energy deposit and the particle positions.
@@ 2.2.2 Modeling the position distributions
Approximately 97% of all possible configurations are fitted, with cluster size from 1 to 8 pixels, and saved in a file to be later used in the reconstruction. The remaining 3% are mostly composed of all possible configurations from 9 to 14 pixels, possible only by exceptionnally large amount of energy deposited by the two particles in the silicon. If such configuration is generated in the Monte-Carlo simulation, it would be discarded and regenerated until the cluster matches a configuration saved in the file.
@@ 3 Reconstruction algorithm for a multi-GeV photon polarimeter
A full simulation-reconstruction code has been written such as all experimental parameters could be tested. The list of parameter includes the cell geometry, the number of cells, the density of the Rohacell-like converter, the photon energy and its polarization. Conversions are generated in the first layer of converter and Geant4 [9] takes over the propagation of the electron/positron pair in the various materials. When crosssing the MAPS, the position of both particles in the middle of the silicon plane are stored in a tree. Secondary particles stemming from the interaction of the lepton pair with the matter are ignored in the present study.
@@ 3.1 Reconstructing the two tracks and the azimuthal angle of the conversion
A Kalman filter is run backwards, i.e. from the last layer to the first layer encountered by the lepton pair. Two track candidates are initiated, one for each measurement in the last layer. The position of the particle is set by the measurement and the tangente of the track angle is initialized with the photon angle (assuming a PbWO4 calorimeter is installed behind the pair polarimeter). The position uncertainty is set to the measurement accuracy (4 m) and a multiple of the mean deviation by Coulomb scattering sets the uncertainty on the track angle.
@@ 4.1 Cell length
Here the cell is simply composed of a 1mm-converter of Rohacell-like material with a density of 40 kg/m3, separated from a MAPS plane by a layer of air as described by Figure 9. Then the intrinsic analyzing power is extracted for various length of the air layer and shown in the right plot of Figure 9.
@@ 4.2 Converter-MAPS distance
AI appears to slowly increase with the Converter-MAPS distance for all three energies, although the distance between MAPS is constant. To understand this conclusion, a closer look must be taken on the computation of the azimuthal angle between the position-based and the direction-based methods introduced in subsection 3.1. In Figure 11, distributions of the direction-based (red) and position-based (blue) reconstructed azimuthal angles are shown as function of the distance between the converter and the MAPS. These distributions are filled on an event-by-event basis, only when the associated uncertainty is found smaller than its alternative.
@@ 4.3 Optimization of the radiation and converter lengths
In this situation, the challenge is to “accurately” derive the depth of the conversion and the associated uncertainty. Using the simulation, the distance between the measurements in the first MAPS is studied as function of the conversion depth. Figure 12 displays the bidimensional histogram representing the conversion depth as function of the distance between the centers of position distributions on the first MAPS layer. A straightforward conclusion is that the conversion is most likely occuring deep in the converter if the distance between the two clusters in the MAPS is large. Consequently a minimal depth for the conversion can be inferred from the distance between the two leptons. As the converter has a finite size, the maximal depth is known as well. When the distance between the leptons is greater than 30m, the distribution of conversion is almost uniform as seen by the projection for a 100 m-distance in Figure 12. Therefore leptons are propagated in the middle of the minimal and maximal depth given the measured distance in the first MAPS, with an uncertainty being the available depth divided by . Two examples are illustrated in Figure 12, where red dots and bars represents the assumed conversion depth and the associated uncertainty for measurements distant by 100 and 250 m on the first MAPS layer.
@@ 4.4 Figure-of-merit for a 1m-long polarimeter
Assuming a maximal length of 1m, the polarimeter can be composed of 8 cells. In the previous studies, the Kalman filter was running over all 16 layers of MAPS for conversion occuring in the first cell. It is important to study the analyzing power as function of the number of MAPS layer downstream of the conversion in order to estimate the polarimeter efficiency. Figure 14 displays the analyzing power as function of the number of MAPS layers provided to the Kalman filter, for photon energies from 1 to 22 GeV.

### v2

@@ Abstract
The polarization of multi-GeV photon is an experimental observable holding many promises in multiple fields. In Astrophysics, such high-energy -rays are emitted by supermassive blackholes, pulsars or heavy-mass binary systems. Accessing their polarization would reveal much about the environment and structure of these astrophysical objects [ilie]. The structure of much smaller objects such as hadrons would greatly benefit from the knowledge of the -rays polarization. Indeed deeply virtual Compton scattering is a process in which a virtual photon inelastically scatters off a quark confined in the proton. The polarization of the final state photon is expected to be a key observable to study linearly polarized gluons in the proton, considered as the intrinsic strong field to form the hadron. Finally the structure of vacuum itself can be probed as well by studying the polarization of multi-GeV photons. Exposed to an intense field being either electromagnetic, gravitational or even strong such as in a quark-gluon Plasma, vacuum may become birefringent which would break Lorentz invariance. The birefringence can be highlighted by the alteration of the photon polarization when passing through the vacuum [ilie, QuantumGravity, Birefringence]. Unfortunately such sources or physics phenomena have low brightness or cross sections. Consequently, to measure the associated linear polarization of a multi-GeV photon with a reasonable statistical accuracy, the polarimeter must have a high Figure-of-Merit.
@@ 1.2 State-of-the-art in -polarimetry via pair production
At low energy, 10-20 MeV, a Time Projection Chamber (TPC) offers a resolution good enough to resolve the lepton pair: The photon converts in the gas and the lepton pair is seen by the ionization of the latter from the conversion vertex. A prototype of such a concept named HARPO [HARPO] was successfully built and caracterized by CEA, CNRS and École Polytechnique. However the maximal photon energy for polarimetry is limited by the TPC spatial resolution and the conversion efficiency is low as the converter is a gas. Now there are pair polarimeters with very thin solid converters (0.2% ) and trackers located downstream of a dipole to resolve the lepton pair [de_Jager_2004]. However the size and the low conversion efficiency makes it unable to study polarization of at most one million uncollimated photons. The most relevant design of a pair polarimeter for high-energy -rays is presented in Eingorn et al. [bogdan] and shown in Figure 1.
@@ 2 A revisited concept of pair polarimeter for DVCS
The design of polarimeter proposed by Eignorn et al. has greatly inspired the subsequent work. For multi-GeV photons, a first limitation comes from the spatial resolution of MSDs which is not good enough unless the length of the cell is dramatically increased. Here we suggest to use monolithic active pixel silicon sensors (MAPS). With pixels being 2525 m2, MAPS detectors achieve 5m-resolution. An example of such detector is the Muon Forward Tracker of ALICE consisting of ALPIDE sensors [MAGER2016434]. Another significant advantage is the thinness of the silicon layer being only 50 m with a detection efficiency above 99%. Including the support layers (Kapton, Alumimium, Carbon,…) the material budget in the active region is approximately 0.25% X/X0. Nowadays an intense R&D effort is still ongoing to reduce even more the material budget of MAPS as well as their power consumption. However the low radiation length of MAPS decreases the efficiency of the polarimeter. To compensate this loss without increasing the number of MAPS layer and consequently the cost of the apparatus, an independent converter must be added to the cell design.
Looking back to Eingorn’s design, a second limitation stems from the thinness of the converter: All the multiple scattering happens in the vicinity of the conversion vertex and the information about the conversion angles are immediately/maximally blurred. Here we suggest to fill the cell with a low-density material as displayed by Figure 2. Unlike a thin converter, the multiple scattering in a low-density material will happen all along the track instead of immediately after the conversion. As the information is recovered from the position measurements in the MAPS, a deviation of the particle is much less significant when it occurs close to the measurement planes than at the conversion vertex. The extended converter could be some high-pressure heavy gas or low-density material such as Rohacell 31/51/71.
@@ 2.1 A realistic MAPS description
First, the single-particle response will be optimized to reproduce the ALPIDE characteristics presented in [MAGER2016434]. Then a study of cluster shapes caused by two close particles going through the MAPS plane is performed to extract as much information as possible to be used at a later stage in a Kalman filter for the polarimeter study.
@@ 2.1.1 MAPS response to single particle
where the definitions of the plasma energy and the parameter can be found in the section entitled passage of particles through matter in the PDG [PDG]. The width of the distribution is . Following this distribution, an energy deposit is randomly generated for each particle in the Toy Monte-Carlo. This energy is then converted into a number of ionized electron/hole pairs placed uniformly along the path of the particle through the fully depleted zone.
@@
where is the hole diffusion constant (=10.86 cm2/s). In order to reproduce as best as possible ALPIDE caracteristics presented in the article [MAGER2016434], the MPV and width of the Landau-Vavilov distribution had to be multiplied by 0.28: For very thin absorber such as 50m-thick silicon sensors, the MPV is known to be smaller than the results from the PDG formula. Displayed by Figure 3, both efficiency and cluster size distributions as function of the threshold are faithfully reproduced whereas the resolution is found 1m better than the measured ones.
In addition to these fundamental properties, the average cluster size as function of the particle location in a pixel are compared in Figure 4. When the particle hits the middle of a pixel, the average cluster size is close to 1. As the particle gets closer to a neighbor pixel either on the same row or column, the average cluster size increases to 2. When getting closer to two neighbor pixels by the diagonale of the pixel, L-shaped 3-pixel clusters become the main pattern which may turn into a squared 4-pixel cluster when the particle hits the corner of a pixel.
@@ 2.1.2 MAPS response to two close particles
First, a single particle is shot at a random position in the middle pixel of a 99-pixel matrix. Then, at a random distance of at most 100 m from the first particle, a second particle is generated. The threshold is set to 90 electrons to consider a pixel being fired. The matrix is then analyzed to determine the number of clusters and their relative position resulting from the convolution between energy deposit and the particle positions.
@@ 2.2.2 Modeling the position distributions
Approximately 97% of all possible configurations are fitted, with cluster size from 1 to 8 pixels, and saved in a file to be later used in the reconstruction. The remaining 3% are mostly composed of all possible configurations from 9 to 14 pixels, possible only by exceptionally large amount of energy deposited by the two particles in the silicon. If such configuration is generated in the Monte-Carlo simulation, it would be discarded and regenerated until the cluster matches a configuration saved in the file.
@@ 3 Reconstruction algorithm for a multi-GeV photon polarimeter
A full simulation-reconstruction code has been written so that all experimental parameters could be tested. The list of parameters includes the cell geometry, the number of cells, the density of the Rohacell-like converter, the photon energy and its polarization. Conversions are generated in the first layer of converter and Geant4 [GEANT4] takes over the propagation of the electron/positron pair in the various materials. When crossing the MAPS, the position of both particles in the middle of the silicon plane are stored in a tree. Secondary particles stemming from the interaction of the lepton pair with the matter are ignored in the present study.
@@ 3.1 Reconstructing the two tracks and the azimuthal angle of the conversion
A Kalman filter is run backwards, i.e. from the last layer to the first layer encountered by the lepton pair. Two track candidates are initiated, one for each measurement in the last layer. The position of the particle is set by the measurement and the tangent of the track angle is initialized with the photon angle (assuming a PbWO4 calorimeter is installed behind the pair polarimeter). The position uncertainty is set to the measurement accuracy (4 m) and a multiple of the mean deviation by Coulomb scattering sets the uncertainty on the track angle.
@@ 4.1 Cell length
Here the cell is simply composed of a 1mm-converter of Rohacell-like material with a density of 40 kg/m3, separated from a MAPS plane by a layer of air as described by Figure 9. Then the intrinsic analyzing power is extracted for various lengths of the air layer and shown in the right plot of Figure 9.
@@ 4.2 Converter-MAPS distance
AI appears to slowly increase with the Converter-MAPS distance for all three energies, although the distance between MAPS is constant. To understand this conclusion, a closer look must be taken on the computation of the azimuthal angle between the position-based and the direction-based methods introduced in subsection 3.1. In Figure 11, distributions of the direction-based (red) and position-based (blue) reconstructed azimuthal angles are shown as a function of the distance between the converter and the MAPS. These distributions are filled on an event-by-event basis, only when the associated uncertainty is found smaller than its alternative.
@@ 4.3 Optimization of the radiation and converter lengths
In this situation, the challenge is to “accurately” derive the depth of the conversion and the associated uncertainty. Using the simulation, the distance between the measurements in the first MAPS is studied as a function of the conversion depth. Figure 12 displays the bidimensional histogram representing the conversion depth as a function of the distance between the centers of position distributions on the first MAPS layer. A straightforward conclusion is that the conversion is most likely occuring deep in the converter if the distance between the two clusters in the MAPS is large. Consequently a minimal depth for the conversion can be inferred from the distance between the two leptons. As the converter has a finite size, the maximal depth is known as well. When the distance between the leptons is greater than 30m, the distribution of conversion is almost uniform as seen by the projection for a 100 m-distance in Figure 12. Therefore leptons are propagated in the middle of the minimal and maximal depth given the measured distance in the first MAPS, with an uncertainty being the available depth divided by . Two examples are illustrated in Figure 12, where red dots and bars represent the assumed conversion depth and the associated uncertainty for measurements distant by 100 and 250 m on the first MAPS layer.
@@ 4.4 Figure-of-merit for a 1m-long polarimeter
Assuming a maximal length of 1m, the polarimeter can be composed of 8 cells. In the previous studies, the Kalman filter was running over all 16 layers of MAPS for conversion occuring in the first cell. It is important to study the analyzing power as a function of the number of MAPS layer downstream of the conversion in order to estimate the polarimeter efficiency. Figure 14 displays the analyzing power as a function of the number of MAPS layers provided to the Kalman filter, for photon energies from 1 to 22 GeV.